# Testing the Segment Anything Model (SAM) on radiology data

### v1

@@ 2.2 Zero-shot segmentation
Zero-shot segmentation involves the generation of masks for new, previously unseen classes [20], with several different methods arising in recent years. Bucher and others use a text model to embed classes which are mapped into a feature space common to both pixel embeddings and text, allowing them to map new words (classes) to pixel embedding space [20]. Björn et. al [21] use a generative model to learn the distribution of the unseen classes, which included a novel decoder architecture that allowed for the generation of class-specific features. This model would then use descriptions of the targets to generate features that would be used by a classifier to perform the segmentation. Luddecke et. al [22] use a dual-encoder CLIP-based [23] architecture to handle both image prompts as well as supporting text prompts that share a common decoder used for the segmentation. Não entendi bem como é queue isto see traduz em zero/one-shot segmentation Bian et. al [11] propose a two-stage approach for multi-modality medical image segmentation, where during the first stage, a fully supervised CNN is trained to extract relation prototypes from a given modality, which are then used to train the zero-shot model for additional modalities.
@@ 3.4 Performance evaluation
Having used either of the four heuristics outlined above, we focus on two relevant metrics: the Dice score ($\mathrm{Dice}=\frac{2\times\mathrm{TP}}{2\times\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}$, where $\mathrm{TP}$, $\mathrm{FP}$ and $\mathrm{FN}$ are the true positives, false positives and false negatives, respectively). To assess whether the model performs well on specific slices, we also calculate the Dice score for each volume considering only slices with at least one valid predicted segment (we henceforth refer to this as $\mathrm{Dice_{pos}}$). Finally, we also calculate the ratio between the number of ground truth pixels in slices where at least one segment of interest was found and the total number of ground truth pixels. We term this last quantity the "fraction of detected object". All metrics are calculated individually for each sample and the metrics distributions for each zone and dataset are analysed.
@@ 4.1 SAM underperforms on radiology data
We applied SAM to 5 different radiology segmentation datasets and calculated the IoU with the ground truth for each individual sample. In general, we observe that SAM does not perform particularly well when segmenting any of the relevant anatomical structures across datasets (Figure 1, left). Indeed, the best performance is observed for the segmentation of the left atrium in heart MRI data ($\mathrm{Dice}=65\%$ when considering masks with > 0.25 IoU) — in other words, if one accepts segments with a relatively small overlap, it is possible to get a perhaps acceptable segmentation with SAM. Considering the best performance for each dataset, SAM performs the worst when detecting colon cancer in colon CT ($\mathrm{Dice}=4.2\%$). We do however admit that this is a particularly challenging segmentation task, with the best performing model of the Medical Segmentation Decathlon achieving a Dice score of 56%, the lowest in the challenge [37].
@@ 4.3 Seeds do not lead to improved performance
To test whether seeds (point prompts for SAM) can ameliorate the poor performance observed thus far, we conducted a simple experiment: using the ProstateX dataset, we take the available whole gland masks and calculate a $3\times 3$ grid of equally spaced points spanning the whole prostate (Figure 4). Using these points, we prompt SAM to obtain prostate gland masks. We observe that this fails to improve Dice scores (Figure 5). In fact, controlling for segment selection heuristic, we observe an unexpected drop of 4% Dice when using seeded SAM (Table 2). This suggests that, for radiology, SAM benefits more from having a more aggressive seeding mechanism (by default SAM uses a $32\times 32$ grid of equally spaced coordinates as seeds) than from having more targeted seeds.
@@ 6 Conclusion
$\approx$25%
$\approx$50%
Best
Figure 6: Examples of SAM’s segmentations of medical images sorted by IoU.

### v2

@@ 2.2 Zero-shot segmentation
Zero-shot segmentation involves the generation of masks for new, previously unseen classes [20], with several different methods arising in recent years. Bucher and others use a text model to embed classes which are mapped into a feature space common to both pixel embeddings and text, allowing them to map new words (classes) to pixel embedding space [20]. Björn et. al [21] use a generative model to learn the distribution of the unseen classes, which included a novel decoder architecture that allowed for the generation of class-specific features. This model would then use descriptions of the targets to generate features that would be used by a classifier to perform the segmentation. Luddecke et. al [22] use a dual-encoder CLIP-based [23] architecture to handle both image prompts as well as supporting text prompts that share a common decoder used for the segmentation. Bian et. al [11] propose a two-stage approach for multi-modality medical image segmentation, where during the first stage, a fully supervised CNN is trained to extract relation prototypes from a given modality, which are then used to train the zero-shot model for additional modalities.
@@ 3.4 Performance evaluation
Having used either of the four heuristics outlined above, we focus on the Dice score ($\mathrm{Dice}=\frac{2\times\mathrm{TP}}{2\times\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}$, where $\mathrm{TP}$, $\mathrm{FP}$ and $\mathrm{FN}$ are the true positives, false positives and false negatives, respectively). To assess whether the model performs well on specific slices, we also calculate the Dice score for each volume considering only slices with at least one valid predicted segment (we henceforth refer to this as $\mathrm{Dice_{pos}}$). Finally, we also calculate the ratio between the number of ground truth pixels in slices where at least one segment of interest was found and the total number of ground truth pixels. We term this last quantity the "fraction of detected object". All metrics are calculated individually for each sample and the metrics distributions for each zone and dataset are analysed.
@@ 4.1 SAM underperforms on radiology data
We applied SAM to 5 different radiology segmentation datasets and calculated the IoU with the ground truth for each individual sample. In general, we observe that SAM does not perform particularly well when segmenting any of the relevant anatomical structures across datasets (Figures 1, left, and 6). Indeed, the best performance is observed for the segmentation of the left atrium in heart MRI data ($\mathrm{Dice}=65\%$ when considering masks with > 0.25 IoU) — in other words, if one accepts segments with a relatively small overlap, it is possible to get a perhaps acceptable segmentation with SAM. Considering the best performance for each dataset, SAM performs the worst when detecting colon cancer in colon CT ($\mathrm{Dice}=4.2\%$). We do however admit that this is a particularly challenging segmentation task, with the best performing model of the Medical Segmentation Decathlon achieving a Dice score of 56%, the lowest in the challenge [37].
@@ 4.3 Seeds do not lead to improved performance
To test whether seeds (point prompts for SAM) can ameliorate the poor performance observed thus far, we conducted a simple experiment: using the ProstateX dataset, we take the available whole gland masks and calculate a $3\times 3$ grid of equally spaced points spanning the whole prostate (Figure 4). Using these points, we prompt SAM to obtain prostate gland masks. We observe that this fails to improve Dice scores (Figure 5). In fact, controlling for segment selection heuristic, we observe an unexpected drop of 4% Dice when using seeded SAM (Table LABEL:tab:linear-seeded). This suggests that, for radiology, SAM benefits more from having a more aggressive seeding mechanism (by default SAM uses a $32\times 32$ grid of equally spaced coordinates as seeds) than from having more targeted seeds.
@@ 6 Conclusion
$\approx$25$\approx$50Best
Figure 6: Examples of SAM’s segmentations of medical images sorted by IoU. The blue outlines represent the ground truth, while the red binary masks represent the outputted segmentation by SAM.